OpenStudy (anonymous):

Calc Question: Force 1 = 70 lbs Force 2 = 100 lbs Resultant = 80 lbs What is the angle between the forces given the resultant force is the magnitude.

OpenStudy (anonymous):

dot product 80 = 70*100*cos(theta) cos(theta) = 8/700 theta is around 89.35 degrees

OpenStudy (anonymous):

I have seen similar problems worked this way, but the book yields the answer as 53 degrees. Any thoughts?

OpenStudy (anonymous):

70^2 + 100^2 + 2(70)(100)cos(theta) = 80^2 14000cos(theta) = -8500 cos(theta) = -17/28 theta = 127.383 adjust for negative 53

OpenStudy (anonymous):

^-- law of cosine

OpenStudy (anonymous):

sorry for mistake :x

OpenStudy (anonymous):

Excuse my unintelligent moment...but may I inquire on how you got the -8500.

OpenStudy (anonymous):

80^2 - 70^2 - 100^2 = -8500 i moved all numbers to right side

OpenStudy (anonymous):

AH! Alright. Thanks for the help *Medal Given*

OpenStudy (anonymous):

lol thanks and no problem